OpenStudy (anonymous):

Find an equation of a hyperbola with a=452 units and c=765 units. Assume that the transverse axis is horizontal. a.x^2/204,304 *y^2/380,921=1 b.x^2/380,921*y^2/204,304=1 c.x^2/452*y^2/313=1 d.x^2/313-y^2/452=1

OpenStudy (dumbcow):

since its horizontal, the a^2 goes under the x^2 to find b^2, use a^2 +b^2 = c^2

OpenStudy (anonymous):

can u explain it, cause my teacher didnt explain it that well

OpenStudy (dumbcow):

not much more to explain for a hyperbola opening in horizontal direction: \[\frac{x^{2}}{a^{2}} - \frac{y^{2}}{b^{2}} = 1\] a^2 +b^2 = c^2 you are given a and c, solve for b

OpenStudy (anonymous):

Would it be C?